Find the equation of the line which passes through the points A (-2, 7) and B (2, -3)


1
Expert's answer
2021-09-06T16:20:28-0400

Slope of the line: "m=\frac{7+3}{-2-2}=-\frac{5}{2}."

Equation of the line: "y-7=-\frac{5}{2}(x+2)" or "y=-\frac{5}{2}x+2."